# Pitfalls of Evidence-Based AI Policy

### v4

@@ 1 How do We Regulate Emerging Tech?
Recently, debates over AI governance have been ongoing across the world. A common underlying theme is the challenge of regulating emerging technologies amidst uncertainty about the future. Even among people who strongly agree that it is important to regulate AI, there is sometimes disagreement about when and how. This uncertainty has led some researchers to call for “evidence-based AI policy.”
@@ 1.2 A Broad, Emerging Coalition
Recently, there have been a number of prominent calls for evidence-based AI policy. For example, several California congressmembers and Governor Gavin Newsom recently argued against an AI regulatory bill in California by highlighting that it was motivated by mitigating future risks that have not been empirically observed:
@@
They follow this with an argument against the precautionary principle (Taleb et al., 2014), claiming that policymakers should take a noncommittal approach in the face of uncertainty and not act on speculative estimates of future AI risks.
@@
Overall, the evidence-based AI policy coalition is diverse. It includes a variety of policymakers and researchers who do not always agree with each other. We caution against developing a one-dimensional view of this coalition or jumping to conclusions from quotes out of context. However, this camp is generally characterized by a desire to avoid pursuing highly committal policy absent compelling evidence.
@@ 1.3 A Vague Agenda?
Calls for evidence-based policy are not always accompanied by substantive recommendations. However, Bommasani et al. (2024a) end their article with a set of four milestones for researchers and policymakers to pursue: [^1] Bommasani et al. (2024a) also call for the establishment of a registry, evaluation, red-teaming, incident reporting, and monitoring but do not specify any particular role for regulators to play in these. They also make a nonspecific call for policymakers to broadly invest in risk analysis research and to investigate transparency requirements.
@@
These milestones are extremely easy to agree with. Unfortunately, they are also unworkably vague. It is unclear what it would mean for them to be accomplished. In fact, for these milestones, it is not hard to argue that existing reports reasonably meet them. For example, the AI Risk Repository (Slattery et al., 2024) predates Bommasani et al. (2024a) and offers a meta-review, taxonomy, and living database of AI risks discussed in the literature. If this does not offer a workable taxonomy of risks (Milestone 1), it is unclear what would. [^2] For milestone 2, most relevant research is domain-specific; consider Metta et al. (2024), Sandbrink (2023), Musser (2023), and Cazzaniga et al. (2024) as examples. Note, however, that forecasting future marginal risks will always be speculative to some degree. See also Bengio et al. (2025). Meanwhile, milestones 3 and 4 essentially describe the first and second stages of the AI regulation process, so existing regulatory efforts already are working on these (e.g., Arda, 2024).
These milestones are an encouraging call to actively improve our understanding. However, absent more precision, we worry that similar arguments could be misused as a form of tokenism to muddy the waters and stymie policy action. In the rest of this paper, we will argue that holding regulatory action to too high an evidentiary standard can paradoxically make it harder to gather the information that we need for good governance.
@@ 2 The Evidence is Biased
In its pure form, science is a neutral process. But it is never done in a vacuum. Beneath the cloak of objectivity, there are subjective human beings working on problems that were not randomly selected (Kuhn, 1962). There is a laundry list of biases subtly shaping the evidence produced by AI researchers. A policymaking approach that fixates on existing evidence to guide decision-making will systematically neglect certain problems.
@@ 2.1 Selective Disclosure
In February 2023, Microsoft announced Bing Chat, an AI-powered web browsing assistant. It was a versatile, semi-autonomous copilot to help users browse the web. It was usually helpful, but sometimes, it went off the rails. Users found that it occasionally took on shockingly angsty, deceptive, and outright aggressive personas. It would go so far as to threaten users chatting with it. Rest assured, everyone was fine. Bing Chat was just a babbling web app that could not directly harm anyone. But it offers a cautionary tale. Right now, developers are racing to create increasingly agentic and advanced AI systems (Chan et al., 2023; Casper et al., 2025). If more powerful future systems go off the rails in similar ways, it could spell trouble.
@@
Historically, AI research and development has been a very open process. For example, code, models, and methodology behind most state-of-the-art AI systems were broadly available pre-2020. More recently, however, developers like Microsoft have been exercising more limited and selective transparency (Bommasani et al., 2024b). Due to a lack of accountability in the tech industry, some lessons remain simply out of reach. There is a mounting crisis of transparency in AI when it is needed the most.
@@
Bengio et al. (2024), International Scientific Report on the Safety of Advanced AI
Differences in the measurability of different phenomena can cause insidious problems to be neglected. For instance, compare explicit and implicit social biases in modern language models. Explicit biases from LLMs are usually easy to spot. For example, it is relatively easy to train a language model against expressing harmful statements about a demographic group. But even when we do this to language models, they still consistently express more subtle biases in the language and concept associations that they use to characterize different people (Wan et al., 2023; Wan & Chang, 2024; Bai et al., 2024; Hofmann et al., 2024).
Meanwhile, benchmarks provide the main scaffolding behind research progress in AI (Patterson, 2012; Hendrycks & Woodside, 2022). For example, tests like GPQA (Rein et al., 2023) actively serve to guide progress on language model capabilities. Many of the benchmarks used in AI research are designed with the hope that they can help us understand downstream societal impacts. However, the strengths of benchmarks are also their weaknesses. Standardized, simplified, and portable measures of system performance often make for poor proxies to study real-world impacts (Raji et al., 2021). For example, in a systematic study of benchmarks designed to assess harmlessness in AI systems, Ren et al. (2024) found that many existing benchmarks intended to evaluate these qualities were, in practice, more reflective of a model’s general capabilities than anything else.
@@
For example, AI ethics researchers have contrasted India and the West to highlight the challenges posed by cultural homogeneity in the research community. In the West, societal discussions around fairness can, of course, be very nuanced, but they reflect Western experiences and are often characterized by a focus on race and gender politics. In India, however, the axes of social disparity are different and, in many ways, more complex. For example, India has 22 official languages, a greater degree of religious conflict, and a historical Caste system. This has led researchers to argue that the AI community is systematically poised to neglect many of the challenges in India and other non-Western parts of the world (Sambasivan et al., 2021; Bhatt et al., 2022; Qadri et al., 2023).
@@ 2.5 The Culture & Values of the AI Research Community
They found an overwhelming predominance of values pertaining to system performance (green) over the other categories of user rights and ethical principles. This suggests that the AI community may be systematically predisposed to produce evidence that will disproportionately highlight the benefits of AI compared to its harms.
@@ 2.6 Industry’s Entanglement with Research
To see an example of Big Tech entangled with calls for “evidence-based AI policy,” we need to look no further than Bommasani et al. (2024a): A Path for Science‑ and Evidence‑based AI Policy (discussed above in Section 1). Five of its seventeen authors have undisclosed for-profit industry affiliations. [^4] Li, Pineau, Varoquaux, Stoica, Liang These include the vice president of AI research at Meta and cofounders of World Labs, Together.ai, Databricks, Anyscale, and :probabl, each of which might be affected by future AI regulations. [^5] The original version of the article did not contain any disclaimers about omitted author affiliations. However, it was updated in October 2024 to disclaim that “Several authors have unlisted affiliations in addition to their listed university affiliation. This piece solely reflects the authors’ personal views and not those of any affiliated organizations.” However, these conflicts of interest are still not disclosed. Failing to disclose clear conflicts of interest in an explicitly political article fails to meet standards for ethical disclosure in research. These standards exist for good reason because a policymaker reading A Path for Science‑ and Evidence‑based AI Policy (Bommasani et al., 2024a) might interpret it very differently if it were clear that some of the authors had obvious conflicts of interest. It is certainly a red flag that calls for more evidence before passing highly committal regulation are coming, in part, from authors with conveniently hidden industry ties.
@@ 3 Lacking Evidence as a Reason to Act
So if the evidence is systematically biased? What do we do? How do we get more, better evidence?
@@
We argue that a need to more thoroughly understand AI risks is a reason to pass regulation–not to delay it. To see this, we first need to understand the distinction between “substantive” regulation and “process” regulation. For our purposes, we define them as such:
@@
These two categories of regulations do not only apply to AI. In the food industry, for example, ingredient bans are substantive regulations while requirements for nutrition facts are process regulations. Process regulations usually pose significantly lower burdens and downsides than substantive ones. The key reason why this distinction is important is that, as we will argue:
@@
3.2 In Defense of Compute & Cost Thresholds in AI Regulation
Some AI policy proposals set cost and compute thresholds such that, if a system’s development surpasses these, it would be subject to specific requirements. Some researchers have rightly pointed out that there are hazards associated with this; cost and compute can be poor proxies for societal risk (Hooker, 2024; Heim & Koessler, 2024).
These are important and needed points about the limitations of cost and compute thresholds. For example, suppose that we are considering substantive regulations that prevent deploying certain models in certain ways. In this case, we would need careful cost-benefit analysis and the ability to adapt regulatory criteria over time. But until we have government agencies who are capable of performing high-quality evaluations of AI systems’ risks, cost and compute thresholds may be the only tenable proxy available (Heim & Koessler, 2024).
In the case of process regulation, there is often a lack of substantial downside. For example, consider policies that require developers to register a system with the government if its development exceeds a cost or compute threshold. Compared to inaction, the upside is a significantly increased ability of the government to monitor the frontier model ecosystem. As for the downside? Sometimes certain companies will accidentally be required to do more paperwork than regulators may have intended. Compared to the laundry list of societal-scale risks from AI (Slattery et al., 2024), we can safely say that this risk is practically negligible.
@@ 4.3 The Duty to Due Diligence from Discoverable Documentation of Dangerous Deeds
Meanwhile, there is legal precedent for companies to sometimes lose court cases when they internally communicate risks through legally discoverable media such as in Grimshaw v. Ford (of Appeal, 1981). So absent requirements, companies will tend to suppress the documentation of dangers to avoid accountability. Meanwhile, mere voluntary transparency can be deceptive by selectively revealing information that reflects positively on the company (Ananny & Crawford, 2018). Thus, we argue that a regime that requires developers to produce scrutable documentation will be key to facilitating the production of more meaningful evidence.
@@ 4.4 Considering Counterarguments
“These 15 objectives would be too burdensome for developers.” It’s true that protocols and documentation can impose burdens. However, these burdens are generally far lighter than those imposed by substantive regulations, and compliance with many of these requirements may be trivial for developers already planning to take similar actions internally. For instance, even the most potentially burdensome measures – such as risk assessments and staged deployments – are practices that major developers like OpenAI, Anthropic, and Google DeepMind have already publicly committed to implementing.
“It’s a slippery slope toward overreach.” A second concern is that these 15 regulatory objectives might generate information that could pave the way for future substantive regulations or liability for developers. Regulatory and liability creep are real phenomena that can harm industry interests. However, it’s important to emphasize that any progression from these objectives to future regulations or liability will ultimately depend on human decision-makers acting on evidence. Evidence is essential for society to engage in meaningful deliberation and exercise informed agency – this is the entire point of evidence-based policy. Therefore, if process-based AI regulations eventually lead to substantive regulations, it won’t be because the process regulations laid an inevitable framework. It would be because the information produced by those regulations persuaded policymakers to take further action. It would be very precarious to argue that a democratic society should be protected from its own access to information about what it is consuming.
@@ 5 Building a Healthier Ecosystem
Process regulations can lay the foundation for more informed debates and decision-making in the future. Right now, the principal objective of AI governance work is not necessarily to get all of the right substantive regulations in place. It is to shape the AI ecosystem to better facilitate the ongoing process of identifying, studying, and deliberating about risks. This requires being critical of the biases shaping the evidence we see and proactively working to seek more information. Kicking the can down the road for a lack of ‘enough’ evidence could impair policymakers’ ability to take needed action.
This lesson is sometimes painfully obvious in retrospect. In the 1960s and 70s, a scientist named S.J. Green was head of research at the British American Tobacco (BAT) company. He helped to orchestrate BAT’s campaign to deny urgency and delay action on public health risks from tobacco. However, he later split with the company, and after reflecting on the intellectual and moral irresponsibility of these efforts, he remarked:

### v6

@@ 1 How do We Regulate Emerging Tech?
Recently, debates over AI governance have been ongoing across the world. A common underlying theme is the challenge of regulating emerging technologies amidst uncertainty about the future. Even among people who strongly agree that it is important to regulate AI, there is sometimes disagreement about when and how. This uncertainty has led some experts and politicians to call for “evidence-based AI policy.”
@@ 1.2 A Broad, Emerging Coalition
Recently, there have been a number of calls for evidence-based AI policy. For example, several California congressmembers and Governor Gavin Newsom recently argued against an AI regulatory bill in California by highlighting that it was motivated by mitigating future risks that have not been empirically observed:
@@
They follow this with an argument against the precautionary principle (Taleb et al., 2014), arguing that speculative estimates of future AI risks are not sufficient to warrant current action.
@@
Overall, the evidence-based AI policy coalition is diverse. It includes a variety of policymakers and researchers who do not always agree with each other. However, this camp is generally characterized by a desire to avoid pursuing highly committal policy absent compelling evidence.
@@ 1.3 A Vague Agenda?
Calls for evidence-based policy are not always accompanied by substantive recommendations. However, here we will discuss one that does. Bommasani et al. (2024a) end their article with a set of four milestones for researchers and policymakers to pursue: [^1] Bommasani et al. (2024a) also call for the establishment of a registry, evaluation, red-teaming, incident reporting, and monitoring but do not specify any particular role for regulators to play in these. They also make a nonspecific call for policymakers to broadly invest in risk analysis research and to investigate transparency requirements.
@@
These milestones are extremely easy to agree with. Unfortunately, they are also likely to be unworkably vague. It is unclear what it would mean for them to be accomplished. On the contrary, it is not hard to argue that existing reports reasonably meet them. For example, the AI Risk Repository (Slattery et al., 2024) predates Bommasani et al. (2024a) and offers a meta-review, taxonomy, and living database of AI risks discussed in the literature. If this does not offer a workable taxonomy of risks (Milestone 1), it seems unclear what would. [^2] For milestone 2, most relevant research is domain-specific; consider Metta et al. (2024), Sandbrink (2023), Musser (2023), and Cazzaniga et al. (2024) as examples. Note, however, that forecasting future marginal risks will always be speculative to some degree. See also Bengio et al. (2025). Meanwhile, milestones 3 and 4 essentially describe the first and second stages of the AI regulation process, so existing regulatory efforts already are working on these (e.g., Arda, 2024).
These milestones are an encouraging call to actively improve our understanding. However, absent more precision, we worry that similar arguments could be misused as a form of tokenism to muddy the waters and stymie policy action. In the rest of this paper, we will argue that holding regulatory action to too high an evidentiary standard can paradoxically make it harder to gather the information that we need for beneficial AI governance.
@@ 2 The Evidence is Biased
In its pure form, science is a neutral process. But it is never done in a vacuum. Beneath the cloak of objectivity, there are subjective human beings working on problems that were not randomly selected (Kuhn, 1962). There is a laundry list of biases subtly shaping the evidence produced by AI researchers. A policymaking approach that fixates on existing evidence to guide decision-making will systematically favor Western, technocratic, industry, and risk-tolerant interests
@@ 2.1 Selective Disclosure
In February 2023, Microsoft announced Bing Chat, an AI-powered web browsing assistant. It was a versatile, semi-autonomous copilot to help users browse the web. It was usually helpful, but sometimes, it went off the rails. Users found that it occasionally took on shockingly angsty, deceptive, and outright aggressive personas. It would go so far as to threaten users chatting with it. It was not a tragic incident – everyone was fine. Bing Chat was just a babbling web app that could not directly harm anyone. But it offers a cautionary tale. Right now, developers are racing to create increasingly agentic and advanced AI systems (Chan et al., 2023; Casper et al., 2025). If more powerful future systems go off the rails in similar ways, it could spell trouble.
@@
The lack of transparency around Bing Chat is no isolated occurrence. Historically, AI research and development has been a very open process. For example, code, models, and methodology behind most state-of-the-art AI systems were broadly available pre-2020. More recently, however, developers like Microsoft have been exercising more limited and selective transparency (Bommasani et al., 2024b). Due to a lack of accountability in the tech industry, some lessons remain simply out of reach. There is a mounting crisis of transparency in AI when it is needed the most.
@@
Bengio et al. (2024), International Scientific Report on the Safety of Advanced AI, Interim Report
Differences in the measurability of different phenomena can cause pernicious problems to be neglected. For instance, compare explicit and implicit social biases in modern language models. Explicit biases from LLMs are usually easy to spot. For example, it is relatively easy to train a language model against expressing harmful statements about a demographic group. But even when we do this to language models, they still consistently express more subtle biases in the language and concept associations that they use to characterize different people (Wan et al., 2023; Wan & Chang, 2024; Bai et al., 2024; Hofmann et al., 2024).
Meanwhile, benchmarks are the backbone of research progress in AI (Patterson, 2012; Hendrycks & Woodside, 2022). For example, tests like GPQA (Rein et al., 2023) actively serve to guide progress on language model capabilities. Many of the benchmarks used in AI research are designed with the hope that they can help us understand downstream societal impacts. However, the strengths of benchmarks are also their weaknesses. Standardized, simplified, and portable measures of system performance often make for poor proxies to study real-world impacts (Raji et al., 2021). For example, in a systematic study of benchmarks designed to assess harmlessness in AI systems, Ren et al. (2024) found that many existing benchmarks intended to evaluate these qualities were, in practice, more reflective of a model’s general capabilities than anything else.
@@
As a specific example, AI ethics researchers often contrast India and the West to highlight the challenges posed by cultural homogeneity in the research community. In the West, societal discussions around fairness can, of course, be very nuanced, but they reflect Western experiences and are often characterized by a focus on race and gender politics. In India, however, the axes of social disparity are different and, in many ways, more complex. For example, India has 22 official languages, a greater degree of religious conflict, and a historical Caste system. This has led researchers to argue that the AI community is systematically poised to neglect many of the challenges in India and other non-Western parts of the world (Sambasivan et al., 2021; Bhatt et al., 2022; Qadri et al., 2023).
@@ 2.5 The Culture & Values of the AI Research Community
There was an overwhelming predominance of values pertaining to system performance (green) over the other categories of user rights and ethical principles. This suggests that the AI community may be systematically predisposed to produce evidence that will disproportionately highlight the benefits of AI compared to its harms.
@@ 2.6 Industry’s Entanglement with Research
To see a potential example of Big Tech entangled with calls for “evidence-based AI policy,” we can scrutinize the authors behind Bommasani et al. (2024a): A Path for Science‑ and Evidence‑based AI Policy (discussed above in Section 1). By our analysis, five of its seventeen authors have undisclosed for-profit industry affiliations. [^4] Li, Pineau, Varoquaux, Stoica, Liang These include the vice president of AI research at Meta and cofounders of World Labs, Together.ai, Databricks, Anyscale, and :probabl, each of which might be affected by future AI regulations. [^5] The original version of the article did not contain any disclaimers about omitted author affiliations. However, it was updated in October 2024 to disclaim that “Several authors have unlisted affiliations in addition to their listed university affiliation. This piece solely reflects the authors’ personal views and not those of any affiliated organizations.” However, these conflicts of interest are still not disclosed. Not disclosing conflicts of interest in an explicitly political article does not meet generally accepted standards for ethical disclosure in research. We argue that in this case in particular, a policymaker reading A Path for Science‑ and Evidence‑based AI Policy (Bommasani et al., 2024a) might interpret it very differently if it were clear that some of the authors had industry conflicts of interest.
@@ 3 Lacking Evidence as a Reason to Act
So if the evidence is systematically biased, what do we do? How do we get more, better evidence?
@@
We argue that the need to more thoroughly understand AI risks is a reason to pass regulation–not to delay it. To see this, we first highlight the distinction between “substantive” regulation and “process” regulation. For our purposes, we define them as such:
@@
These two categories do not only apply to AI. In the food industry, for example, ingredient bans are substantive regulations while requirements for nutrition facts are process regulations. Process regulations usually pose significantly lower burdens and downsides than substantive ones. The key reason why this distinction is important is that, as we will argue:
@@
3.2 Defense of Compute & Cost Thresholds in AI Regulation
Some AI policy proposals set cost and compute thresholds such that, if a system’s development surpasses these, it would be subject to specific requirements. Some researchers have duly pointed out that there are hazards associated with this; cost and compute can be poor proxies for societal risk (Hooker, 2024; Heim & Koessler, 2024).
These are important and needed points about the limitations of cost and compute thresholds. For example, suppose that we are considering substantive regulations that prevent deploying certain models in certain ways. In this case, we would need careful cost-benefit analysis and the ability to adapt regulatory criteria over time. But until we have government agencies who are capable of performing high-quality evaluations of AI systems’ risks, thresholds may be the only tenable proxy available (Heim & Koessler, 2024).
In the case of process regulation, there is often a lack of substantial downside. For example, consider policies that require developers to register a system with the government if its development exceeds a cost or compute threshold. Compared to inaction, the upside is a significantly increased ability of the government to monitor the frontier model ecosystem. As for the downside? Sometimes, certain companies will accidentally be required to do more paperwork and be more transparent than regulators may have intended. Compared to the laundry list of societal-scale risks from AI (Slattery et al., 2024), we argue that this type of risk is practically negligible.
@@ 4.3 The Duty to Due Diligence from Discoverable Documentation of Dangerous Deeds
Meanwhile, there is legal precedent for companies to sometimes lose court cases when they internally communicate risks through legally discoverable media such as in Grimshaw v. Ford (of Appeal, 1981). So absent requirements, companies will tend to actively suppress the documentation of dangers to avoid accountability. Meanwhile, mere voluntary transparency can be deceptive by selectively revealing information that reflects positively on the company (Ananny & Crawford, 2018). Thus, we argue that a regime that requires developers to produce scrutable documentation will be key to facilitating the production of more meaningful evidence.
@@ 4.4 Considering Counterarguments
“These 15 objectives would be too burdensome for developers.” It is true that protocols and documentation can impose burdens. However, these burdens are generally far lighter than those imposed by substantive regulations, and compliance with many of these requirements may be trivial for developers already planning to take similar actions internally. For instance, even the most potentially burdensome measures – such as risk assessments and staged deployments – are practices that major developers like OpenAI, Anthropic, and Google DeepMind have already publicly committed to implementing.
“It’s a slippery slope toward overreach.” A second concern is that these 15 regulatory objectives might generate information that could pave the way for future substantive regulations or liability for developers. Regulatory and liability creep are real phenomena that can harm industry interests. However, it is important to emphasize that any progression from these objectives to future regulations or liability will ultimately depend on human decision-makers acting on evidence. Evidence is essential for society to engage in meaningful deliberation and exercise informed agency – this is the entire point of evidence-based policy. Therefore, if process-based AI regulations eventually lead to substantive regulations, it will not be because the process regulations laid an inevitable framework. It would be because the information produced by those regulations persuaded policymakers to take further action. Ultimately, we believe it would be very precarious to argue that a democratic society should be protected from its own access to information about what it is consuming.
@@ 5 Building a Healthier Ecosystem
Process regulations can lay the foundation for more informed debates and decision-making in the future. Right now, the principal objective of AI governance work is not necessarily to get all of the right substantive regulations in place. Instead, we argue that it is to shape the AI ecosystem to help us better identify, study, and deliberate about risks. This requires being critical of the biases shaping the evidence we see and proactively working to seek more information. Kicking the can down the road for a lack of ‘enough’ evidence could impair policymakers’ ability to take needed action.
This lesson is sometimes painfully obvious in retrospect. For example, in the 1960s and 70s, a scientist named S.J. Green was head of research at the British American Tobacco (BAT) company. He helped to orchestrate BAT’s campaign to deny urgency and delay action on public health risks from tobacco. However, he later split with the company, and after reflecting on the intellectual and moral irresponsibility of these efforts, he remarked: